# Leveraging Joint Predictive Embedding and Bayesian Inference in Graph Self Supervised Learning

### v1

@@ Abstract
Graph representation learning has emerged as a cornerstone for tasks like node classification and link prediction, yet prevailing self-supervised learning (SSL) methods face challenges such as computational inefficiency, reliance on contrastive objectives, and representation collapse. Existing approaches often depend on feature reconstruction, negative sampling, or complex decoders, which introduce training overhead and hinder generalization. Further, current techniques which address such limitations fail to account for the contribution of node embeddings to a certain prediction in the absence of labeled nodes. To address these limitations, we propose a novel joint embedding predictive framework for graph SSL that eliminates contrastive objectives and negative sampling while preserving semantic and structural information. Additionally, we introduce a semantic-aware objective term that incorporates pseudo-labels derived from Gaussian Mixture Models (GMMs), enhancing node discriminability by evaluating latent feature contributions. Extensive experiments demonstrate that our framework outperforms state-of-the-art graph SSL methods across benchmarks, achieving superior performance without contrastive loss or complex decoders. Key innovations include (1) a non-contrastive, view-invariant joint embedding predictive architecture, (2) Leveraging single context and multiple targets relationship between subgraphs, and (3) GMM-based pseudo-label scoring to capture semantic contributions. This work advances graph SSL by offering a computationally efficient, collapse-resistant paradigm that bridges spatial and semantic graph features for downstream tasks. The code for our paper can be found at https://github.com/Deceptrax123/JPEB-GSSL
1 Introduction
@@
Graph representation learning has found widespread adoption in Social Network Analysis, Recommendation Systems, Computer Vision and Natural Language Processing. It aims to learn low-dimensional embeddings of nodes or subgraphs while preserving their underlying spatial and spectral features. These learned embeddings can then be used on downstream tasks such as node classification([1]), link prediction([2]) and community detection([3]) by training task-specific decoders or classification layers on the learned embeddings keeping the backbone frozen. Such an approach reduces the computational complexity and training time on downstream tasks. Though Graph Neural Networks(GNNs) have gained popularity over time, they require a certain number of labelled nodes to perform well. Further, owing to the complexity of graph representations spatially, it is often challenging to pre-train graphs and transfer the learned embeddings to downstream tasks.
@@
Graph Self Supervised learning has been widely studied in the literature to facilitate graph representation learning. This includes methods such as Node2Vec, DGI([4]), MVGRL([5]), GRACE([6]) etc. Although these methods have achieved state-of-the-art results on several social networks, they rely heavily on either graph reconstruction/feature reconstruction, re-masking or generating contrastive views by negative sampling, which is a computationally expensive process on large graphs. Graph-SSL methods that rely on reconstruction (generative methods) are decoder variant, meaning they depend on the decoder architecture(inner product decoder or symmetrical feature reconstruction). Such networks might require additional feature propagation steps such as skip connections to prevent vanishing gradients and retain learned information. Furthermore, most methods tend to use more layers or stack multiple encoders to capture long-range node interactions, which increases training complexity and leads to poor performance in downstream tasks due to over-smoothing. This diminishes the model’s ability to effectively discriminate between nodes effectively([7]).
@@
  • We conduct extensive experiments to show that our approach significantly outperforms previous state-of-the-art Graph Self Supervised Learning(G-SSL) counterparts.
2 Related Work
2.1 Unsupervised Representation Learning on Graphs
@@ 2.2 Bootstrapping Methods
2.3 Joint Predictive Embedding Methods
@@ 3.1 Preliminaries
3.2 Loss Function
3.2.1 Joint Predictive Optimization
@@
$\mathcal{L}^{J}(\Theta^{C},\Theta^{R})=\frac{1}{T}\Sigma_{k=1}^{k=T}||H^{\prime}_{k}-H^{t}_{k}||^{2}$ (1)
where $T$ is the total number of targets, $\Theta^{C}$ is the weight matrix of the context encoder and $\Theta^{R}$ is the weight matrix of the target encoder. The objective is computed at the latent space with no reconstruction/negative sampling involved.
3.2.2 Node Feature Contribution Optimization
The node embeddings $H^{\prime}$ obtained from passing $G^{\prime}$ into the context encoder are fit in a Gaussian Mixture Model(GMM). We aim to obtain,
$p(z_{k}=1|h^{\prime}_{n})=\frac{p(h^{\prime}_{n}|z_{k}=1)p(z_{k}=1)}{\Sigma_{j=1}^{j=K}p(h^{\prime}_{n}|z_{j}=1)p(z_{j}=1)}$ (2)
where $z$ is a latent variable that takes 2 values: one if $h^{\prime}$ comes from Gaussian $k$, and zero otherwise. We can obtain,
@@ 3.2.2 Node Feature Contribution Optimization
where $\pi_{k}$ is the mixing coefficient such that,
$\Sigma_{k=1}^{k=K}\pi_{k}=1$ (5)
On replacing these equations in eq.2, we obtain
$p(z_{k}=1|h^{\prime}_{n})=\frac{\pi_{k}\mathcal{N}(h^{\prime}_{n}|\mu_{k},\Sigma_{k})}{\Sigma\pi_{j}\mathcal{N}(h^{\prime}_{n}|\mu_{j},\Sigma_{j})}=\gamma(z_{nk})$ (6)
E-Step: In the E-Step, we aim to evaluate,
$\mathcal{Q}(\theta^{*},\theta)=\mathbb{E}[\ln p(H^{\prime},Z|\theta^{*})]=\Sigma_{Z}p(Z|H^{\prime},\theta)\ln p(H^{\prime},Z|\theta^{*})$ (7)
from eq.6, we can substitute $\gamma$ in the above equation as,
$\mathcal{Q}(\theta^{*},\theta)=\Sigma_{Z}\gamma(z_{nk})\ln p(H^{\prime},Z|\theta^{*})$ (8)
On finding the complete likelihood of the model, we finally have,
$\mathcal{Q}(\theta^{*},\theta)=\Sigma_{n=1}^{N}\Sigma_{k=1}^{K}\gamma(z_{nk})[\ln\pi_{k}+\ln\mathcal{N}(h^{\prime}_{n}|\mu_{k},\Sigma_{k})]$ (9)
M-Step: In the M-Step, we aim to find updated parameters $\theta^{*}$ as follows,
@@
Considering the restriction, $\Sigma_{k=1}^{K}\pi_{k}=1$, eq.9 is modified as follows,
$\mathcal{Q}(\theta^{*},\theta)=\Sigma_{n=1}^{N}\Sigma_{k=1}^{K}\gamma(z_{nk})[\ln\pi_{k}+\ln\mathcal{N}(h^{\prime}_{n}|\mu_{k},\Sigma_{k})]-\lambda(\Sigma_{k=1}^{k=K}\pi_{k}-1)$ (11)
The parameters are then determined by finding the maximum likelihood of $\mathcal{Q}$. On taking the derivative with respect to $\pi_{k},\mu_{k},\Sigma_{k}$ and rearranging terms, we finally obtain the update equations for the parameters as follows,
$\pi_{k}=\frac{\Sigma_{n=1}^{n=N}\gamma(z_{nk})}{N}$ (12)
@@
$\sum_{k}^{*}=\frac{\sum_{n=1}^{N}\gamma(z_{nk})(h^{\prime}_{n}-\mu_{k})(h^{\prime}_{n}-\mu_{k})^{T}}{\sum_{n=1}^{N}\gamma(z_{nk})}$ (14)
@@
$\mathcal{L}^{G}=\begin{cases}0.5(\frac{V_{g}^{T}H^{\prime}}{\Sigma V_{g}||V_{g}||}-\frac{V_{k}^{T}H^{\prime}}{\Sigma V_{k}||V_{k}||})/\beta,&\text{if}|\frac{V_{g}^{T}H^{\prime}}{\Sigma V_{g}||V_{g}||}-\frac{V_{k}^{T}H^{\prime}}{\Sigma V_{k}||V_{k}||}|<\beta\\ |\frac{V_{g}^{T}H^{\prime}}{\Sigma V_{g}||V_{g}||}-\frac{V_{k}^{T}H^{\prime}}{\Sigma V_{k}||V_{k}||}|-0.5\beta,&\text{otherwise}\par\end{cases}$ (15)
@@ 3.2.3 Final objective
3.3 Description of Components and Operations
3.3.1 Context Encoder
@@ 3.3.3 Predictor
3.3.4 Global Mean Pooling Operation
@@
where $x_{n}$ refers to the node features of node $n$ and $N$ is the total number of nodes in graph $G$
4 Experiments
4.1 Experimental Setting
4.1.1 Datasets
In our experiments, we evaluate our proposed framework on six publicly available benchmark datasets for node representation learning and classification. The datasets are, namely, Cora([17]), Pubmed([18]), Citeseer([17]), Amazon Photos([19]), Amazon Computers([19]) and Coauthor CS([19]). Cora consists of 2708 scientific publications classified into one of seven classes, Citeseer consists of 3312 scientific publications classified into one of six classes and Pubmed consists of 19717 scientific publications pertaining to diabetes classified into one of three classes. In Amazon Photos and Amazon Computers, nodes represent goods and edges represent that two goods are frequently bought together. The product reviews are represented as bag-of-words features. Coauthor CS contains paper keywords for each author’s papers. Nodes represent authors that are connected by an edge if they co-authored a paper.
4.1.2 Evaluation Methodology
The embeddings are evaluated on node classification, whose performance is quantified by accuracy. We follow the same evaluation protocol as used in [20], [12], [11] etc.
4.1.3 Evaluation Protocol
For all downstream tests, we follow the linear evaluation protocol on graphs where the parameters of the backbone/encoder are frozen during inference time. Only the prediction head, which is a single GCN layer, is trained for node classification. For evaluation purposes, we use the default number for train/val/test splits for the citation networks i.e. Cora, Pubmed, Citeseer which are 500 validation and 1000 testing nodes. The train/val/test splits for the remaining datasets, namely, Amazon Photos, Amazon Computers and Coauthor CS are according to [19]. Unless otherwise mentioned, performance is averaged over 10 independent runs with different random seeds and splits for all six evaluation datasets. We report the mean and standard deviation obtained across 10 runs.
4.1.4 Environment
@@
4.2 Evaluation Results
4.2.1 Performance on Node Classification
As mentioned earlier, we use the linear evaluation protocol and report the mean and standard deviation of classification accuracy on the test nodes over 10 runs on different folds or splits with different seeds. We compare our proposed approach with supervised and fine-tuned model baselines. For supervised node classification, we compare our proposed approach against Multi Layer Perceptron(MLP),Graph Convolution Network(GCN)([21]), Graph Attention Network(GAT)([20]), Simplified GCN([22]), Logistic Regression and GraphSAGE([23]). For GraphSAGE, we use the mean, maxpool and meanpool variants as described in [19]. We have compared our model’s performance against supervised baselines in table 1. For self-supervised and fine-tuned node classification, we compare our proposed approach against DGI, MVGRL([5]), GRACE([6]), CCA-SSG([9]), SUGRL([24]), S3-CL([12]), GraphMAE([10]), GMI([25]), BGRL([13]) and ParetoGNN([11]). Table 2 contains our model’s performance against self-supervised baselines. The proposed approach demonstrates superior performance across both small and large datasets, showcasing its scalability and robustness. In supervised node classification (Table 1), our method achieves the highest accuracy on Cora ($89.8\pm 0.9$), Citeseer ($77.0\pm 0.9$), and Amazon Photos ($94.5\pm 0.5$), outperforming state-of-the-art models like BGRL and GraphMAE. Specifically, compared to BGRL, our approach shows significant improvements on Cora and Citeseer, while maintaining competitive results on Pubmed and Coauthor CS. Table 3 highlights a direct comparison with BGRL and ParetoGNN, where our method outperforms both on Photos ($94.5\pm 0.5$) and CS ($93.6\pm 0.4$), and remains competitive on Computers. These results underscore the approach’s adaptability and strong performance across diverse dataset sizes, particularly excelling over BGRL and ParetoGNN in key benchmarks. Owing to the ability of our proposed approach to avoid noisy features, representation collapse, and leverage semantic information, our proposed model performs well on small and unstable datasets such as cora and citeseer and large datasets such as Amazon and Coauthor.
Method Cora Citeseer Pubmed Amazon P Amazon Comp. Coauthor CS

### v2

@@ Abstract
Graph representation learning has emerged as a cornerstone for tasks like node classification and link prediction, yet prevailing self-supervised learning (SSL) methods face challenges such as computational inefficiency, reliance on contrastive objectives, and representation collapse. Existing approaches often depend on feature reconstruction, negative sampling, or complex decoders, which introduce training overhead and hinder generalization. Further, current techniques which address such limitations fail to account for the contribution of node embeddings to a certain prediction in the absence of labeled nodes. To address these limitations, we propose a novel joint embedding predictive framework for graph SSL that eliminates contrastive objectives and negative sampling while preserving semantic and structural information. Additionally, we introduce a semantic-aware objective term that incorporates pseudo-labels derived from Gaussian Mixture Models (GMMs), enhancing node discriminability by evaluating latent feature contributions. Extensive experiments demonstrate that our framework outperforms state-of-the-art graph SSL methods across benchmarks, achieving superior performance without contrastive loss or complex decoders. Key innovations include (1) a non-contrastive, view-invariant joint embedding predictive architecture, (2) Leveraging single context and multiple targets relationship between subgraphs, and (3) GMM-based pseudo-label scoring to capture semantic contributions. This work advances graph SSL by offering a computationally efficient, collapse-resistant paradigm that bridges spatial and semantic graph features for downstream tasks.
Index Terms:
graph neural networks, node classification, self-supervised learning
I Introduction
@@
Graph representation learning has found widespread adoption in Social Network Analysis, Recommendation Systems, Computer Vision and Natural Language Processing[1]. It aims to learn low-dimensional embeddings of nodes or subgraphs while preserving their underlying spatial and spectral features. These learned embeddings can then be used on downstream tasks such as node classification[2], link prediction[3] and community detection[4] by training task-specific decoders or classification layers on the learned embeddings keeping the backbone frozen. Such an approach reduces the computational complexity and training time on downstream tasks. Though Graph Neural Networks(GNNs) have gained popularity over time, they require a certain number of labelled nodes to perform well. Further, owing to the complexity of graph representations spatially, it is often challenging to pre-train graphs and transfer the learned embeddings to downstream tasks.
@@
Graph Self Supervised learning has been widely studied in the literature to facilitate graph representation learning. This includes methods such as Node2Vec, DGI[5], MVGRL[6], GRACE[7] etc. Although these methods have achieved state-of-the-art results on several social networks, they rely heavily on either graph reconstruction/feature reconstruction, re-masking or generating contrastive views by negative sampling, which is a computationally expensive process on large graphs. Graph-SSL methods that rely on reconstruction (generative methods) are decoder variant, meaning they depend on the decoder architecture(inner product decoder or symmetrical feature reconstruction). Such networks might require additional feature propagation steps such as skip connections to prevent vanishing gradients and retain learned information. Furthermore, most methods tend to use more layers or stack multiple encoders to capture long-range node interactions, which increases training complexity and leads to poor performance in downstream tasks due to over-smoothing. This diminishes the model’s ability to effectively discriminate between nodes effectively[8].
@@
  • We conduct extensive experiments demonstrating that our approach significantly outperforms previous state-of-the-art Graph Self-Supervised Learning (G-SSL) methods. Additionally, we provide a thorough emperical efficiency analysis highlighting the model’s efficiency and scalability.
II Related Work
II-A Unsupervised Representation Learning on Graphs
@@ II-B Bootstrapping Methods
II-C Joint Predictive Embedding Methods
@@ III-A Preliminaries
III-B Loss Function
III-B1 Joint Predictive Optimization
@@
$\mathcal{L}^{J}(\Theta^{C},\Theta^{R})$ $=$ $\frac{1}{T}\sum_{k=1}^{T}\|H^{\prime}_{k}-H^{t}_{k}\|^{2}$ (1)
where $T$ is the total number of targets, $\Theta^{C}$ is the weight matrix of the context encoder, and $\Theta^{R}$ is the weight matrix of the target encoder. The objective is computed at the latent space with no reconstruction/negative sampling involved.
III-B2 Node Feature Contribution Optimization
The node embeddings $H^{\prime}$ obtained from passing $G^{\prime}$ into the context encoder are fit in a Gaussian Mixture Model (GMM). We aim to obtain:
$p(z_{k}=1|h^{\prime}_{n})$ $=$ $\frac{p(h^{\prime}_{n}|z_{k}=1)p(z_{k}=1)}{\sum_{j=1}^{K}p(h^{\prime}_{n}|z_{j}=1)p(z_{j}=1)}$ (2)
where $z$ is a latent variable that takes two values: one if $h^{\prime}$ comes from Gaussian $k$, and zero otherwise. We can obtain:
@@ III-B2 Node Feature Contribution Optimization
where $\pi_{k}$ is the mixing coefficient such that:
$\sum_{k=1}^{K}\pi_{k}$ $=$ 11 11 (5)
By substituting these equations into (2), we obtain:
$p(z_{k}=1|h^{\prime}_{n})$ $=$ $\frac{\pi_{k}\mathcal{N}(h^{\prime}_{n}|\mu_{k},\Sigma_{k})}{\sum_{j=1}^{K}\pi_{j}\mathcal{N}(h^{\prime}_{n}|\mu_{j},\Sigma_{j})}=\gamma(z_{nk})$ (6)
E-Step: In the E-Step, we aim to evaluate:
$\mathcal{Q}(\theta^{*},\theta)$ $=$ $\mathbb{E}[\ln p(H^{\prime},Z|\theta^{*})]$ (7)
$=$ $\sum_{Z}p(Z|H^{\prime},\theta)\ln p(H^{\prime},Z|\theta^{*})$
From (6), we substitute $\gamma$ into the equation:
$\mathcal{Q}(\theta^{*},\theta)$ $=$ $\sum_{Z}\gamma(z_{nk})\ln p(H^{\prime},Z|\theta^{*})$ (8)
The complete likelihood of the model then becomes:
$\mathcal{Q}(\theta^{*},\theta)$ $=$ $\sum_{n=1}^{N}\sum_{k=1}^{K}\gamma(z_{nk})\left[\ln\pi_{k}+\ln\mathcal{N}(h^{\prime}_{n}|\mu_{k},\Sigma_{k})\right]$ (9)
M-Step: In the M-Step, we aim to find updated parameters $\theta^{*}$ as follows:
@@
Considering the constraint $\sum_{k=1}^{K}\pi_{k}=1$, equation (9) is modified as follows:
$\mathcal{Q}(\theta^{*},\theta)$ $=$ $\sum_{n=1}^{N}\sum_{k=1}^{K}\gamma(z_{nk})\Big{[}\ln\pi_{k}+\ln\mathcal{N}(h^{\prime}_{n}|\mu_{k},\Sigma_{k})\Big{]}$ (11)
$-\lambda\left(\sum_{k=1}^{K}\pi_{k}-1\right)$
The parameters are then determined by maximizing $\mathcal{Q}$. Taking the derivative with respect to $\pi_{k},\mu_{k},\Sigma_{k}$ and rearranging terms, we obtain the update equations:
$\pi_{k}$ $=$ $\frac{\sum_{n=1}^{N}\gamma(z_{nk})}{N}$ (12)
@@
$\Sigma_{k}^{*}$ $=$ $\frac{\sum_{n=1}^{N}\gamma(z_{nk})(h^{\prime}_{n}-\mu_{k})(h^{\prime}_{n}-\mu_{k})^{T}}{\sum_{n=1}^{N}\gamma(z_{nk})}$ (14)
@@
$\mathcal{L}^{G}$ $=$ $\begin{cases}\frac{1}{2}\left(\dfrac{V_{g}^{T}H^{\prime}}{\sum V_{g}\|V_{g}\|}-\dfrac{V_{k}^{T}H^{\prime}}{\sum V_{k}\|V_{k}\|}\right)/\beta,\\ \quad\text{if}\ \left|\dfrac{V_{g}^{T}H^{\prime}}{\sum V_{g}\|V_{g}\|}-\dfrac{V_{k}^{T}H^{\prime}}{\sum V_{k}\|V_{k}\|}\right|<\beta,\\[10.0pt] \left|\dfrac{V_{g}^{T}H^{\prime}}{\sum V_{g}\|V_{g}\|}-\dfrac{V_{k}^{T}H^{\prime}}{\sum V_{k}\|V_{k}\|}\right|-\frac{1}{2}\beta,\\ \quad\text{otherwise}.\end{cases}$ (15)
@@ III-B3 Final objective
III-C Description of Components and Operations
III-C1 Context Encoder
@@ III-C3 Predictor
III-C4 Global Mean Pooling Operation
@@
where $x_{n}$ refers to the node features of node $n$ and $N$ is the total number of nodes in graph $G$.
III-D Implementation Details
For all experiments, we use the Adam Optimizer[18] and cosine annealing learning rate scheduler with warm restarts. No regularization techniques such as dropout, $L_{1}$ or $L_{2}$ have been employed since they tend to reduce performance. For learning rates, we perform a search with values corresponding to the search space $\{0.5,0.1,0.05,0.01,0.001\}$. We set the value of the momentum update parameter $m$ to 0.9. The number of epochs was set to 50,000 with early stopping. All experiments were conducted on an M2 Macbook Pro with 8GB CPU. We did not utilize GPUs for any of our experiments.
IV Experiments
IV-A Experimental Setting
IV-A1 Datasets
In our experiments, we evaluate our proposed framework on seven publicly available benchmark datasets for node representation learning and classification. The datasets are, namely, Cora[19], Pubmed[20], Citeseer[19], Amazon Photos[21], Amazon Computers[21], Coauthor CS[21] and WikiCS[22]. Details on number of nodes, edges and features are given in table I Cora consists of 2708 scientific publications classified into one of seven classes, Citeseer consists of 3312 scientific publications classified into one of six classes and Pubmed consists of 19717 scientific publications pertaining to diabetes classified into one of three classes. In Amazon Photos and Amazon Computers, nodes represent goods and edges represent that two goods are frequently bought together. The product reviews are represented as bag-of-words features. Coauthor CS contains paper keywords for each author’s papers. Nodes represent authors that are connected by an edge if they co-authored a paper. WikiCS is a dataset derived from Wikipedia with nodes corresponding to Computer Science articles and edges based on hyperlinks. The 10 classes represent different branches of Computer Science.
TABLE I: Dataset Details
Dataset Nodes Edges Features Classes
Cora 2,708 10,556 1,433 7
Citeseer 3,327 9,104 3,703 6
Pubmed 19,717 88,648 500 3
Amazon Photos 7,650 238,612 745 8
Amazon Computers 13,752 491,722 767 10
Coauthor CS 18,333 163,788 6,805 15
WikiCS 11,701 216,123 300 10
IV-A2 Evaluation Methodology
The embeddings are evaluated on node classification, whose performance is quantified by accuracy. We follow the same evaluation protocol as used in [23], [13], [12] etc. All scores for baselines have been obtained from previously published papers.
IV-A3 Evaluation Protocol
For all downstream tests, we follow the linear evaluation protocol on graphs where the parameters of the backbone/encoder are frozen during inference time. Only the prediction head, which is a single GCN layer, is trained for node classification. For evaluation purposes, we use the default number for train/val/test splits for the citation networks i.e. Cora, Pubmed, Citeseer which are 500 validation and 1000 testing nodes. The train/val/test splits for the remaining datasets, namely, Amazon Photos, Amazon Computers and Coauthor CS are according to [21]. For WikiCS we use the publicily available splits. Unless otherwise mentioned, performance is averaged over 10 independent runs with different random seeds and splits for all seven evaluation datasets. We report the mean and standard deviation obtained across 10 runs.
IV-A4 Environment
@@
IV-B Evaluation Results
IV-B1 Performance on Node Classification on small and large graphs
As mentioned earlier, we use the linear evaluation protocol and report the mean and standard deviation of classification accuracy on the test nodes over 10 runs on different folds or splits with different seeds. We compare our proposed approach with supervised and fine-tuned model baselines. For semi-supervised node classification, we compare our proposed approach against Multi Layer Perceptron(MLP),Graph Convolution Network(GCN)[24], Graph Attention Network(GAT)[23], Simplified GCN[25], Logistic Regression and GraphSAGE[26]. For GraphSAGE, we use the mean, maxpool and meanpool variants as described in [21]. We have compared our model’s performance against semi-supervised baselines in table II. For self-supervised and fine-tuned node classification, we compare our proposed approach against DGI, MVGRL[6], GRACE[7], CCA-SSG[10], SUGRL[27], S3-CL[13], GraphMAE[11], GMI[28], BGRL[14] and ParetoGNN[12] on small and large graphs. In semi-supervised node classification (Table II), our method achieves the highest accuracy across all baselines. On self-supervised classification followed by fine-tuning, we evaluate our proposed model against two specific fields of data i.e Planetoid datasets and larger, more stable datasets such as Amazon Computers, Photos and WikiCS. Table III compares our model’s performance against strong baselines on Planetoid datasets. We show that our model performs extremely well on smaller datasets by outperforming all baselines on Cora and Citeseer with extremely competitive results on Pubmed. Table IV contains the performance of our model against strong baselines for Photos, Computers, Coauthor CS and WikiCS. Our model consistently achieves competitive scores against baselines by outperforming baselines on Photos and CS while remaining competitive on Computers and WikiCS .These results underscore the approach’s adaptability and strong performance across diverse dataset sizes. Owing to the ability of our proposed approach to avoid noisy features, representation collapse, and leverage semantic information, our proposed model performs well on small and unstable datasets such as Cora and Citeseer and large datasets such as Amazon, Coauthor and WikiCS.
TABLE II: Semi-Supervised Node Classification. The values indicate the accuracy achieved on node classification by several methods. A higher value indicates better performance. The best score is marked in bold and the second-best score is underlined. N/A indicates that the score was not reported for a particular dataset by the original authors.
Method Cora Citeseer Pubmed Photos Computers Coauthor CS